Two circles with radii r1 and r2 touch externally. What is the length of their direct common tangent?

Grade:11

3 Answers

Jitender K Yadav
7 years ago
Just draw two circles with centers C and C`, which touch each other at P. (Draw Figure yourself as per instructions.)The direct common tangent PP` has length `d`.CC`=r1+r2Draw PP1||CC` with P1 on C`P`.CP||C`P` (tangent common)Thus CPP1C` is a ||gm.Thus, PP1=CC`=r1+r2P`P1=C`P`-P1P`=r2-r1Since, PP` is perpendicular to P`P1,d=โˆš[(r1+r2)^2-(r2-r1)^2]=โˆš[4r1r2]=2โˆš(r1r2)
gurram meghana
7 years ago
The formula for the above question is ย 2(r1)(r2)/ r1+r2. Therefore the solution for the above Question is ย  2(1)(2)/1+2=4/3
ย 
gurram meghana
7 years ago
The formula to the above question is 2(r1)(r2)/r1+r2. Therefore the solution to the above question is 2(1)(2)/1+2=4/3.